# Zeroth-order log-concave sampling

### v1

@@ 5 Log-concave sampling under an evaluation oracle
To extend our approach to log-concave distributions, we generalize what we did for uniform sampling, streamlining algorithms in [KV25a] with their samplers replaced by our new samplers.
@@
We begin with a cleaner analysis for log-concave sampling from a warm start in §5.1 and then move onto an order-preserving annealing algorithm for warm-start generation in §5.2. To this end, we improve the complexity of a sampler for these annealing distributions.
5.1 Zeroth-order log-concave sampling with balanced guarantees
We take the exponential-lifting approach ($\mathsf{exp\text{-}lifting}$) studied in [KV25b], improving their sampler and simplifying analysis similarly to the uniform sampling case.
@@ 5.1 Zeroth-order log-concave sampling with balanced guarantees
Since the convergence rate of $\mathsf{PS}$ was already analyzed in §2.1, we can just reuse Proposition 2.4. The only adjustment is that we now require a bound on $C_{\mathsf{PI}}(\pi^{X,T})$, rather than $C_{\mathsf{PI}}(\pi^{X})$, as we run $\mathsf{PS}$ to sample from $\pi^{X,T}$. Since $C_{\mathsf{PI}}(\pi^{X,T})\lesssim\lVert\operatorname{cov}\pi^{X,T}\rVert\log d$ due to log-concavity of $\pi^{X,T}$, and in [KV25b, Lemma 2.5], $\lVert\operatorname{cov}\pi^{X,T}\rVert$ is bounded as
@@
Per-step query complexity.
Denoting by $\ell(y)=\mathbb{P}_{\mathcal{N}(y-h\alpha,hI_{d+1})}(\cdot\in\mathcal{K})$ the success probability of rejection sampling, one can check (or see [KV25b, Lemma 2.11]) that for $z=(x,t)\in\mathbb{R}^{d}\times\mathbb{R}$,
$(\pi^{Z}*\gamma_{h})(y)=\pi^{Y}(y)=\frac{\ell(y)\exp(-\alpha^{\mathsf{T}}y+\frac{1}{2}\,h\,|\alpha|^{2})}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}z)\,\mathrm{d}z}\,.$
This has a concentration phenomenon: for $\widetilde{\mathcal{K}}:=\mathcal{K}_{\delta}+h\alpha$, if $\delta\geq 13hd$,
$\pi^{Y}(\widetilde{\mathcal{K}}^{c})\leq\exp\bigl{(}-\frac{\delta^{2}}{2h}+13\delta d\bigr{)}\,.$
Hereafter, we take $h=(2\cdot 13^{2}cd^{2})^{-1}$ and $\delta=(13d)^{-1}$ for some $c>0$, under which $\delta\geq 13hd$ is equivalent to $c\geq 1/2$. Recall the following expansion lemma from [KV25b, Lemma 2.13]:
$\frac{\int_{\mathcal{K}_{s}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}\leq e^{13sd}\quad\text{for }s>0\,.$
This leads to the following bound: using change of variables via $y=v+h\alpha$ in $(i)$,
$\frac{\int_{\widetilde{\mathcal{K}}}\exp(-\alpha^{\mathsf{T}}y+\frac{1}{2}\,h\,|\alpha|^{2})\,\mathrm{d}y}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}\underset{(i)}{=}\frac{\int_{\mathcal{K}_{\delta}}\exp(-\alpha^{\mathsf{T}}v-\frac{1}{2}\,h\,|\alpha|^{2})\,\mathrm{d}v}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}\leq\frac{\int_{\mathcal{K}_{\delta}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}\leq e^{13\delta d}=e\,.$ (5.1)
As in the analysis for $\mathsf{PS}_{\textup{unif}}$, we show that the failure probability of a stage in the $k$-th iteration is at most $1/2$ under suitable choices of $h$ and $\tau$. Let $(p,q)$ be a pair of conjugates with $p,q\geq 2$. Then, $\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]\leq M_{q}\,\{\mathbb{E}_{\pi^{Y}}[(1-\ell)^{\tau p}]\}^{1/p}$ and
$\int_{\mathbb{R}^{d+1}}(1-\ell)^{\tau p}\,\mathrm{d}\pi^{Y}=\int_{\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}}}\cdot+\int_{\widetilde{\mathcal{K}}\cap[\ell\tau\geq\log(3M_{q})]}\cdot+\int_{\widetilde{\mathcal{K}}\cap[\ell\tau<\log(3M_{q})]}\cdot=:\mathsf{A}+\mathsf{B}+\mathsf{C}\,.$
Under $c\asymp\log(3M_{q})^{p}$ and $\tau\asymp M_{q}^{p}\log(3M_{q})$, the last term can be bounded by $(\nicefrac{{1}}{{2M_{q}}})^{p}$ due to
$\mathsf{A}$ $\leq\pi^{Y}(\mathbb{R}^{d+1}\backslash\widetilde{\mathcal{K}})\leq e^{-c+1}\,,$
$\mathsf{B}$ $\leq\int_{\widetilde{\mathcal{K}}\cap[\ell\tau\geq\log(3M_{q})]}\exp(-\ell\tau p)\,\mathrm{d}\pi^{Y}\leq\bigl{(}\frac{1}{3M_{q}}\bigr{)}^{p}\,,$
$\mathsf{C}$ $\leq\int_{\widetilde{\mathcal{K}}\cap[\ell\tau<\log(3M_{q})]}\frac{\ell(y)\exp(-\alpha^{\mathsf{T}}y+\frac{1}{2}\,h\,|\alpha|^{2})}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}\,\mathrm{d}y\leq\frac{\log(3M_{q})}{\tau}\,\frac{\int_{\widetilde{\mathcal{K}}}\exp(-\alpha^{\mathsf{T}}y+\frac{1}{2}\,h\,|\alpha|^{2})\,\mathrm{d}y}{\int_{\mathcal{K}}\exp(-\alpha^{\mathsf{T}}v)\,\mathrm{d}v}$
(5.1)${\leq}\tau^{-1}e\log(3M_{q})\,.$ ) .
Hence, $\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]\leq 1/2$, and the expected number of trials during the $k$-th iteration can be bounded by $2\tau$ just as in §2.2. Putting all these together, we can establish the query complexity of $\mathsf{PS}_{\textup{exp}}$ with balanced guarantees.
@@
For $M:=\mathcal{R}_{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})$, we set $h^{-1}\asymp d^{2}M$ and $\tau\asymp Me^{M}$. As discussed earlier, $\mathsf{PS}_{\textup{exp}}$ suffices to iterate at most
$N\lesssim qMd^{2}(\Lambda\vee 1)\log d\log\frac{M}{\varepsilon}\quad\text{times}\,.$
Thus, the expected number of total queries during $N$ iterations is
$2\tau N\lesssim qM^{2}e^{M}d^{2}(\Lambda\vee 1)\log d\log\frac{M}{\varepsilon}=\widetilde{O}\bigl{(}qd^{2}\,(\Lambda\vee 1)e^{\mathcal{R}_{q}(\pi_{0}^{X}\mathbin{\|}\pi^{X})}\log\frac{1}{\varepsilon}\bigr{)}\,,$
which completes the proof. ∎
5.2 Rényi-warmness generation for log-concave distributions
[…]
@@ 5.2 Rényi-warmness generation for log-concave distributions
which satisfies $\pi(\mathbb{R}^{d+1}\backslash\bar{\mathcal{K}})\leq 1/2$ (see [KV25b, Convex truncation and §3.1]). Thus, we can ensure that $\bar{\pi}=\pi|_{\bar{\mathcal{K}}}$ is $O(1)$-close to $\pi$ in $\mathcal{R}_{\infty}$. Below, we let $D:=Rl$ and $\mu_{i}:=\mu_{\sigma_{i}^{2},\rho_{i}}$, and denote by $\bar{\mu}_{i}$ the law of an actual sample such that $\mathcal{R}_{3}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$.
@@ 5.2.1 Proximal sampler for tilted Gaussian
We now improve a sampler for the annealing distributions $\mu_{i}$ proposed in [KV25b]. For $v:=(x,t)\in\mathbb{R}^{d}\times\mathbb{R}$ and $w:=(y,s)\in\mathbb{R}^{d}\times\mathbb{R}$, consider the augmented target given by
@@
  • [Backward] $v_{k+1}\sim\mu^{V|W=w_{k+1}}\propto\bigl{[}\mathcal{N}(ry_{k+1},h_{r}I_{d})\otimes\mathcal{N}(s_{k+1}-\rho h,h)\bigr{]}\big{|}_{\bar{\mathcal{K}}}$. To this end, draw $v_{k+1}\sim\mathcal{N}(ry_{k+1},h_{r}I_{d})\otimes\mathcal{N}(s_{k+1}-\rho h,h)$ until $v_{k+1}\in\bar{\mathcal{K}}$. If this rejection loop exceeds $\tau$ trials, discard $w_{k+1}$ and return to the forward step using the same $V=v_{k}$.
We establish its query complexity as follows:

### v2

@@ 5 Logconcave sampling under an evaluation oracle
To extend our approach to logconcave distributions, we generalize what we did for uniform sampling, streamlining algorithms in [KV25a] and obtaining a guarantee for $\mathcal{R}_{q}$-warmness generation.
@@
We begin with analysis for order-preserving logconcave sampling from an $O(1)$-warm start in §5.1 and move to an order-preserving annealing algorithm for warm-start generation in §5.2. Overall, we follow a similar route for the analysis as in the uniform sampling case.
5.1 Zeroth-order logconcave sampling with balanced guarantees
We use the exponential-lifting approach ($\mathsf{exp\text{-}lifting}$) introduced in [KV25b], obtaining an order-preserving guarantee under an $O(1)$-warm start. For $z=(x,t)\in\mathbb{R}^{d+1}$ and $y\in\mathbb{R}^{d+1}$, $\mathsf{PS}_{\textup{exp}}$ alternates $y\sim\pi^{Y|Z=z}=\mathcal{N}(z,hI_{d+1})$ and $z\sim\pi^{Z|Y=y}\propto\mathcal{N}(y-h\alpha,hI_{d+1})|_{\mathcal{K}}$ for $\alpha=de_{d+1}$, where the second step is implemented via rejection sampling, using the proposal $\mathcal{N}(y-h\alpha,hI_{d+1})$ with threshold $\tau$.
@@ 5.1 Zeroth-order logconcave sampling with balanced guarantees
Since the convergence rate of $\mathsf{PS}$ was already given in §2.1, we can just reuse Proposition 2.4. The only adjustment is that we now require a bound on $C_{\mathsf{PI}}(\pi^{X,T})$, rather than $C_{\mathsf{PI}}(\pi^{X})$, as we run $\mathsf{PS}$ to sample from $\pi^{X,T}$. Since $C_{\mathsf{PI}}(\pi^{X,T})\lesssim\lVert\operatorname{cov}\pi^{X,T}\rVert\log d$ due to log-concavity of $\pi^{X,T}$, and in [KV25b, Lemma 2.5], $\lVert\operatorname{cov}\pi^{X,T}\rVert$ is bounded as
@@
Per-step analysis.
It follows from §5.1 in [KV25a] that with the parameter choices $S=\frac{16NM_{2}}{\eta}$, $c=\frac{(\log\log S)^{2}}{13^{2}\log S}$, $\tau=Z^{2}\log^{2}Z$, and $h=c/(13d)^{2}$, the failure probability satisfies $\mathbb{E}_{\pi_{k}^{Y}}[(1-\ell)^{\tau}]\leq\frac{\eta}{N}$. The per-step complexity is bounded as follows: setting $q=1+\alpha$ with $\alpha\geq 1\vee\log\tau$, it follows that $\mathbb{E}_{\pi_{k}^{Y}}[\frac{1}{\ell}\wedge\tau]\leq 5M_{q}\log S$.
Proposition 5.1.
For a convex function $V:\mathbb{R}^{d}\to\mathbb{R}$, let $\pi\propto e^{-V}$ be the logconcave distribution over $\mathbb{R}^{d}$ with $B_{1}\subset\mathsf{L}_{\pi,g}$ and $\Lambda=\lVert\operatorname{cov}\pi\rVert$, given access to an evaluation oracle for $V$. Given $\varepsilon>0$, $\eta\in(0,1/2)$, $q\geq 2$, and an initial distribution $\pi_{0}$ with $M_{q}=\lVert\nicefrac{{\mathrm{d}\pi_{0}}}{{\mathrm{d}\pi}}\rVert_{L^{q}(\pi)}$, initialize $\mathsf{PS}_{\textup{exp}}$ with $\pi_{0}$ and iterate it $N=\widetilde{O}(qd^{2}\,(1\vee\Lambda)\log\frac{1}{\varepsilon}\log\frac{1}{\eta})$ times with $h=(13d^{2}\log S)^{-1}$ and $\tau=S^{2}\log^{2}S$ for $S=\frac{16NM_{2}}{\eta}$. If $q\geq 1+\log\tau$ and $M_{q}\leq 10$, then
  • The sufficient condition on $q$ is satisfied when $q\geq 2\vee O(\log\frac{N}{\eta})$.
  • With probability at least $1-\eta$, $\mathsf{PS}_{\textup{exp}}$ iterates $N$ times without failure. Conditioned on this success, the $X$-output $X^{*}\sim\nu$ satisfies $\mathcal{R}_{q}(\nu\mathbin{\|}\pi)\leq\varepsilon+2\log\frac{1}{1-\eta}$, with $\widetilde{O}(qd^{2}\,(1\vee\Lambda)\log\frac{1}{\varepsilon}\log^{2}\frac{1}{\eta})$ evaluation queries in expectation.
Its proof is almost similar to the analysis of $\mathsf{PS}_{\textup{unif}}$ in Proposition 2.8. Putting all these together, we can establish the query complexity of $\mathsf{PS}_{\textup{exp}}$ with balanced guarantees.
@@
Take $\varepsilon=\eta<0.01$, and restart $\mathsf{PS}_{\textup{exp}}$ upon failure. Note that the expected number of restarts is at most $0.02$. In order to ensure $q\geq 2\vee O(\log N)$ for $N=\widetilde{O}(qd^{2}\,(1\vee\Lambda)\log^{2}\frac{1}{\varepsilon})$, it suffices to restrict
$q\geq 2\vee\widetilde{O}\bigl(\log(d^{2}\,(1\vee\Lambda))+\log\log\frac{1}{\varepsilon}\bigr)\,.$
By the proposition above, the desired claims follow in a similar way to the proof of Theorem 1.1. ∎
5.2 Rényi-warmness generation for logconcave distributions
We streamline a warm-start algorithm in [KV25a], which generates a warm start for the lifted distribution $\pi\propto e^{-dt}|_{\mathcal{K}}$ by annealing through distributions $\mu_{\sigma^{2},\rho}(x,t)\propto\exp(-\frac{1}{2\sigma^{2}}\,|x|^{2}-\rho t)\,\mathds{1}_{\bar{\mathcal{K}}}(x,t)$ for some $\bar{\mathcal{K}}$. In comparison with our warm-start algorithm for uniform distributions, we can also stop the $\sigma^{2}$-annealing earlier than [KV25a], but we could not remove the preprocessing step therein, in order to ensure a finite LSI constant of the annealing distributions. Also, since the complexity of sampling from $\mu_{\sigma^{2},\rho}$ (by $\mathsf{PS}_{\textup{ann}}$ in §5.2.1) is roughly $d^{2}(\sigma^{2}\vee 1)$ (from an $O(1)$-warm start), which has the additive $d^{2}$ queries, we cannot help but following a convoluted approach in [KV25a].
@@ 5.2 Rényi-warmness generation for logconcave distributions
which satisfies $\pi(\mathbb{R}^{d+1}\backslash\bar{\mathcal{K}})\leq 1/2$ (see [KV25b, Convex truncation and §3.1]). Thus, we can ensure that $\bar{\pi}=\pi|_{\bar{\mathcal{K}}}$ is $\log 2$-close to $\pi$ in $\mathcal{R}_{\infty}$. Below, we let $D:=Rl$ and $\mu_{i}:=\mu_{\sigma_{i}^{2},\rho_{i}}$, and denote by $\bar{\mu}_{i}$ the law of an actual sample such that $\mathcal{R}_{\widetilde{O}(1)}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$.
@@ 5.2.1 Proximal sampler for tilted Gaussian
Now consider the sampler for the annealing distributions $\mu_{i}$ proposed in [KV25b]. For $v:=(x,t)\in\mathbb{R}^{d}\times\mathbb{R}$ and $w:=(y,s)\in\mathbb{R}^{d}\times\mathbb{R}$, consider the augmented target given by
@@
  • [Backward] $v_{k+1}\sim\mu^{V|W=w_{k+1}}\propto\bigl[\mathcal{N}(ry_{k+1},h_{r}I_{d})\otimes\mathcal{N}(s_{k+1}-\rho h,h)\bigr]\big|_{\bar{\mathcal{K}}}$. To this end, draw $v_{k+1}\sim\mathcal{N}(ry_{k+1},h_{r}I_{d})\otimes\mathcal{N}(s_{k+1}-\rho h,h)$ until $v_{k+1}\in\bar{\mathcal{K}}$. If this rejection loop exceeds $\tau$ trials, declare failure.
We establish its query complexity similarly as before.
Proposition 5.2.
Consider the annealing distribution $\mu=\mu_{\sigma^{2},\rho}$ above, given access to an evaluation oracle for $V$. Given $\varepsilon>0$, $\eta\in(0,1/2)$, $q\geq q_{0}\geq 2$, and an initial distribution $\mu_{0}$ with $M_{q_{0}}=\lVert\nicefrac{{\mathrm{d}\mu_{0}}}{{\mathrm{d}\mu}}\rVert_{L^{q_{0}}(\mu)}$, initialize $\mathsf{PS}_{\textup{ann}}$ with $\mu_{0}$ and iterate it $N=\widetilde{O}(d^{2}\,(\sigma^{2}\vee 1)\log\frac{q}{\varepsilon}\log\frac{1}{\eta})$ times with $h=(24^{2}d^{2}\log S)^{-1}$ and $\tau=2S^{2}\log^{2}S$ for $S=\frac{16NM_{2}}{\eta}$. If $q_{0}\geq 1+\log\tau$ and $M_{q_{0}}\leq 10$, then
  • The sufficient condition on $q_{0}$ is satisfied when $q_{0}\geq 2\vee O(\log\frac{N}{\eta})$.
  • With probability at least $1-\eta$, $\mathsf{PS}_{\textup{ann}}$ iterates $N$ times without failure. Conditioned on this success, the output $Z\sim\nu$ satisfies $\mathcal{R}_{q}(\nu\mathbin{\|}\mu)\leq\varepsilon+2\log\frac{1}{1-\eta}$, with $\widetilde{O}(d^{2}\,(\sigma^{2}\vee 1)\log\frac{q}{\varepsilon}\log^{2}\frac{1}{\eta})$ evaluation queries in expectation.
Proof.
Overall, we follow the analysis of $\mathsf{PS}_{\textup{Gauss}}$ as in Proposition 4.1. Using the Bakry–Émery criterion and bounded perturbation, one can show that $C_{\mathsf{LSI}}(\mu)\lesssim\sigma^{2}\vee 1$ (see [KV25b, Lemma 3.3]). For $q\geq q_{0}>1$, when $\mathsf{PS}_{\textup{ann}}$ starts from $\mu_{0}^{V}$, by Corollary 1.3, in order to be $\varepsilon$-close to $\mu^{V}$ in $\mathcal{R}_{q}$, for $q_{0}>1$, $\mathsf{PS}_{\textup{ann}}$ suffices to iterate
$h^{-1}\,(\sigma^{2}\vee 1)\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\mu_{0}^{V}\mathbin{\|}\mu^{V})}{(2\wedge q_{0}-1)\,\varepsilon}\qquad\text{iterations}\,.$
Also, [KV25a] showed that under $c=\frac{(\log\log S)^{2}}{4\cdot 24^{2}\log S}$, $h=\frac{c}{24^{2}d^{2}}$, and $\tau=2S^{2}\log^{2}S$ for $S=\frac{16NM_{2}}{\eta}$, the per-step query complexity is bounded by $O(M_{q}\log S)$, and the failure probability is bounded by $\eta/N$.
For $q\geq q_{0}\geq 2$ and $(M_{2}\leq)M_{q_{0}}\leq 10$, the required number $N$ of iterations (of $\mathsf{PS}_{\textup{ann}}$) for $\varepsilon$-closeness in $\mathcal{R}_{q}$ should satisfy $N\gtrsim h^{-1}\,(\sigma^{2}\vee 1)\log\frac{q}{\varepsilon}$. Thus, it suffices to take
$N=\widetilde{O}\bigl(d^{2}\,(\sigma^{2}\vee 1)\log\frac{q}{\varepsilon}\log\frac{1}{\eta}\bigr)\,.$
Also, a sufficient condition for ensuring $q_{0}\geq 2\vee\log\tau$ is $q_{0}\geq 2\vee O(\log\frac{N}{\eta})$. Under this condition, the per-step complexity is $O(\log S)$, so the expected number of queries for $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{q}$ is
$\widetilde{O}\bigl(d^{2}\,(\sigma^{2}\vee 1)\log\frac{q}{\varepsilon}\log^{2}\frac{1}{\eta}\bigr)\,$,
and the success probability is at least $1-\eta$. ∎
5.2.2 Query complexity of Rényi-warmness generation
Overall, we follow the same analysis of the annealing for uniform distributions; see ‘High-level idea of analysis’ in §4.3.
Algorithm.
Given $\varepsilon_{\exp}=\eta_{\exp}\in(0,1/10)$ and $q\geq 2$, we pick a base order $q_{0}(\leq q)$ that will be preserved through annealing, and restart the following algorithm from scratch if $\mathsf{PS}_{\textup{ann}}$ declares failure in the middle.
Recall $D=Rl$ (for a large constant $l>0$ such that $\mathcal{R}_{\infty}(\bar{\pi}\mathbin{\|}\pi)\leq 2$) and $\mu_{i}=\mu_{\sigma_{i}^{2},\rho_{i}}$, and $\bar{\mu}_{i}$ is the law of an actual sample such that $\mathcal{R}_{2q_{0}}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})\leq\frac{1}{10}$. For $m$ the number of total phases, $\mathsf{PS}_{\textup{ann}}$ is iterated $N_{i}$ times with $\varepsilon_{i}=1/10$ and $\eta_{i}=(10m)^{-1}$ in each phase $i\in[m]$.
  1. 1.
  2. 2.
    [Phase I: $\sigma^{2}$-warming] While $d^{-1}\leq\sigma_{i}^{2}\leq 1$, move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{1}{(4q_{0}d)^{1/2}}\bigr)\,.$
    At the end of Phase I, run $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$ toward $\mu_{i+1}\propto e^{-|x|^{2}/2-t}|_{\bar{\mathcal{K}}}$, obtaining $\bar{\mu}_{i+1}$.
  3. 3.
    [Phase II: $\rho$-annealing] While $1\leq\rho_{i}\leq d$ (and $\sigma_{i}^{2}\approx 1$), move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{1}{(4q_{0}d)^{1/2}}\bigr)^{-1}\quad\&\quad\rho_{i+1}=\min\bigl(d,\rho_{i}\,\bigl(1+\frac{1}{(4q_{0}d)^{1/2}}\bigr)\bigr)\,.$
    This is followed by inner annealing: until $\sigma_{i+1}^{2}\leq 1$, run $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$ toward new $\mu_{i+1}$ defined by
    $\sigma_{i+1}^{2}\leftarrow\sigma_{i+1}^{2}\,\bigl(1+\frac{\sigma_{i+1}}{q_{0}D}\bigr)\,$,
  4. 4.
    [Phase III: $\sigma^{2}$-annealing] While $\sigma_{i}^{2}\lesssim qD\,(\Lambda^{1/2}\vee 1)\log^{1/2}d$ (and $\rho_{i}=d$), move from $\bar{\mu}_{i}$ to $\bar{\mu}_{i+1}$ by $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{2q_{0}}$, where
    $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,\bigl(1+\frac{\sigma_{i}}{q_{0}D}\bigr)\,.$
    When i=m1i=m-1 (i.e., the last phase), we move from $\bar{\mu}_{m-1}$ to $\bar{\mu}_{m}$ by $\mathsf{PS}_{\textup{ann}}:\mathcal{R}_{q_{0}}\to\mathcal{R}_{q}$.
Choice of parameters.
Note that the number $m$ of inner phases is bounded as
$m\lesssim(q_{0}d)^{1/2}\log d+(q_{0}d)^{1/2}\log d\times\frac{q_{0}^{1/2}D}{d^{1/2}}+q_{0}D\log\bigl(qD\,(\Lambda^{1/2}\vee 1)\bigr)=\widetilde{O}\bigl(q_{0}d^{1/2}D\log q(1\vee\Lambda^{1/2})\bigr)\,.$
Hence, $N_{i}=\widetilde{O}(d^{2}\,(\sigma_{i}^{2}\vee 1)\log^{2}q_{0}m)\leq\widetilde{O}(qd^{2}D\,(\Lambda^{1/2}\vee 1))=:N_{\max}$ and $N_{\exp}:=\widetilde{O}(qd^{2}\,(\Lambda\vee 1)\log\frac{1}{\eta_{\exp}}\log\frac{1}{\varepsilon_{\exp}})$.
We now find sufficient conditions for $q_{0}$ and $q$ to invoke the established sampling guarantees of $\mathsf{PS}_{\textup{ann}}$ and $\mathsf{PS}_{\textup{exp}}$ (see Proposition 5.1 and 5.2); namely, $q\geq 2\vee O(\log\frac{N_{\exp}}{\eta_{\exp}})$, $q_{0}\geq 2\vee O(\log mN_{\max})$, and $q\geq q_{0}$.
Repeating a similar computation as in the uniform-sampling case, we will take $q_{0}=2\vee\widetilde{O}(\log\{qdD\,(\Lambda\vee 1)\})$ and
$q\gtrsim 2\vee\widetilde{O}\bigl(\log\frac{d^{2}\,(\Lambda\vee 1)}{\eta_{\exp}}+\log\log\frac{1}{\varepsilon_{\exp}}\bigr)\vee\widetilde{O}\bigl(\log\{dD\,(\Lambda\vee 1)\}\bigr)\,.$ (5.2)
Also, due to the choice of $\eta_{i}$, the probability that the failure condition is met until the termination of the annealing is bounded by $\frac{1}{10}$.
Complexity analysis.
We first recall a result on closeness of two logconcave distributions from [KV25a, Lemma 4.4].